OpenStudy (anonymous):

write the expression in lowest terms y^2+4y+4/y^2+6y+8

OpenStudy (anonymous):

The best thing you can do here is factor both the numerator and denominator, then you see that a term or two may divide out

OpenStudy (anonymous):

y^2+4y+4=(y+2)(y+2) y^2+6y+8=(y+2)(y+4)

OpenStudy (anonymous):

Therefore: